# On the discontinuities of Hausdorff dimension in generic dynamical Lagrange spectrum

### v1

@@ 3.5. Sequences of subhorseshoes
In this subsection, we suppose existence of an infinite sequence of discontinuities of the map $L_{\varphi,f}$ in some closed sub interval of $I_{\varphi,f}$ that doesn’t contain the first accumulation point of the Lagrange spectrum and then construct arbitrary large finite sequences of subhorseshoes with some specific properties. Observe that here is the first time when we use the hypothesis of the diffeomorphism being close to a conservative one.
Remember that any subhorseshoe $\tilde{\Lambda}_{0}$ of $\Lambda_{0}$ has a continuation $\tilde{\Lambda}\subset\Lambda$ for any $\varphi\in\mathcal{U}$. In theorem $A$ of [15], the authors showed that the maps $D_{\Lambda_{0},u}:\mathcal{U}\rightarrow\mathbb{R}$ and $D_{\Lambda_{0},s}:\mathcal{U}\rightarrow\mathbb{R}$ given by $D_{\Lambda_{0},u}(\varphi)=D_{u}(\Lambda)$ and $D_{\Lambda_{0},s}(\varphi)=D_{u}(\Lambda)$ are continuous and, in fact, the same proof also shows that the continuity of the maps $D_{\tilde{\Lambda}_{0},u}(\varphi)=D_{u}(\tilde{\Lambda})$ and $D_{\tilde{\Lambda}_{0},s}(\varphi)=D_{u}(\tilde{\Lambda})$ is uniform on the subhorseshoes. Moreover, as for $\varphi_{0}$ one can take $\tilde{C}=1$ in 2.4 (see remark 2.2 in [1]) then $D_{u}(\tilde{\Lambda}_{0})=D_{s}(\tilde{\Lambda}_{0})$ for any subhorseshoe $\tilde{\Lambda}_{0}$ of $\Lambda_{0}$ and, as a consequence, we can choose the neighborhood $\mathcal{U}$ of $\varphi_{0}$ small enough such that for some constants $r_{1},r_{2}$ with $r_{1}/r_{2}>999/1000$ and for any subhorseshoe $\tilde{\Lambda}$ of $\Lambda$ one has
(3.11) $r_{1}D_{s}(\tilde{\Lambda})\leq D_{u}(\tilde{\Lambda})\leq r_{2}D_{s}(\tilde{\Lambda}).$
Fix $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$, $f\in\mathcal{P}_{\varphi,\Lambda}$, some closed sub interval $I\subset I_{\varphi,f}$ that doesn’t contain $c_{\varphi,f}$ and suppose we have an infinite sequence of discontinuities of $L_{\varphi,f}$ with $s\in I$ for every $s$ in the sequence. Then, as $L_{\varphi,f}(\min I)\leq L_{\varphi,f}(s)\leq HD(\Lambda_{s})$, by 2.6 and 2.7
@@
Now, in order to get the sequences of subhorseshoes, we will associate to every $n$ a pair of subhorseshoes of $\Lambda$. In fact, the two subhorseshoes $\Lambda^{s}(t_{n})$ and $\Lambda^{u}(t_{n})$ are given by proposition 3.12 considering some $0<\epsilon_{n}<\min\{0.001,(t_{n+1}-t_{n})/2\}$ and they satisfy
@@
  • $\Lambda^{s}(t_{n})$ doesn’t connect with $\Lambda^{u}(t_{n})$ before $t_{n}$,
  • $\Lambda^{s}(t_{n})$ connects with $\Lambda^{u}(t_{n})$ before $t_{n+1}$
@@
We can take $\theta\in\{s,u\}$ such that given $N\in\mathbb{N}$ arbitrary, there exists a sequence $n_{1}<n_{2}<...<n_{N}$ of elements of $\mathcal{I}$ (where $\mathcal{I}$ is the index set of the sequence $\{t_{n}\}$) such that for $i,j\in\{1,...,N\}$ with $i\neq j$$\Lambda^{\theta}(t_{n_{i}})$ and $\Lambda^{\theta}(t_{n_{j}})$ doesn’t connect before $\max\{t_{n_{i}},t_{n_{j}}\}.$
@@
We do the proof when the index set of the sequence is =$\mathcal{I}=\{n\in\mathbb{Z}:n\geq 1\}$, and the other case follows similarly.
In the first case take a maximal $r_{1}$-chain beginning with 1; that is, a $r_{1}$-chain $1=n_{1}<n_{2}<...<n_{r_{1}}$ such that for any $n>n_{r_{1}}$$1=n_{1}<n_{2}<...<n_{r_{1}}<n$ is not a $(r_{1}+1)$-chain and then $\Lambda^{s}(t_{n_{r_{1}}})$ doesn’t connect with $\Lambda^{s}(t_{n})$ before $t_{n}$. Next take a maximal $r_{2}$-chain beginning with $n_{r_{1}}+1$$n_{r_{1}}+1=n^{(r_{1})}_{1}<n^{(r_{1})}_{2}<\dots<n^{(r_{1})}_{r_{2}}$ then, as before, for $n^{(r_{1})}_{r_{2}}<n$$\Lambda^{s}(t_{n^{(r_{1})}_{r_{2}}})$ doesn’t connect with $\Lambda^{s}(t_{n})$ before $t_{n}$. Now consider a maximal $r_{3}$-chain beginning with $n^{(r_{1})}_{r_{2}}+1$:  $n^{(r_{1})}_{r_{2}}+1=n^{(r_{1},r_{2})}_{1}<n^{(r_{1},r_{2})}_{2}<\dots<n^{(r_{1},r_{2})}_{r_{3}}$ then for $n^{(r_{1},r_{2})}_{r_{3}}<n$,  $\Lambda^{s}(t_{n^{(r_{1},r_{2})}_{r_{3}}})$ doesn’t connect with $\Lambda^{s}(t_{n})$ before $t_{n}$.
@@
such that for $k_{1},k_{2}\geq 2$ with $k_{1}\neq k_{2}$,  $\Lambda^{s}(t_{\tilde{n}_{k_{1}}})$ and $\Lambda^{s}(t_{\tilde{n}_{k_{2}}})$ doesn’t connect before $\max\{t_{\tilde{n}_{k_{1}}},t_{\tilde{n}_{k_{2}}}\}$.
On the other hand, in the second case, take $r\in\mathbb{N}$ arbitrarily big and $n_{1}<n_{2}<...<n_{r}$ some $r$-chain, then we affirm that for $i,j\in\{1,...,r\}$ with $i\neq j$,  $\Lambda^{u}(t_{n_{i}})$ and $\Lambda^{u}(t_{n_{j}})$ doesn’t connect before $\max\{t_{n_{i}},t_{n_{j}}\}.$ In other case if for some $i_{0},j_{0}\in\{1,...,r\}$ with $i_{0}<j_{0}$, $\Lambda^{u}(t_{n_{i_{0}}})$ and $\Lambda^{u}(t_{n_{j_{0}}})$ connect before $t_{n_{j_{0}}}$ then as by corollary 2.7,  $\Lambda^{s}(t_{n_{j_{0}}})$ connect with $\Lambda^{s}(t_{n_{i_{0}}})$ before $t_{n_{j_{0}}}$ and as also $\Lambda^{s}(t_{n_{i_{0}}})$ connects with $\Lambda^{u}(t_{n_{i_{0}}})$ before $t_{n_{i_{0}}+1}$ (and then before $t_{n_{j_{0}}}$). Applying two times more that corollary we have that $\Lambda^{s}(t_{n_{j_{0}}})$ connect with $\Lambda^{u}(t_{n_{j_{0}}})$ before $t_{n_{j_{0}}}$ that is a contradiction. From this follows the result. ∎
Without loss of generality, we will suppose that in the previous proposition $\theta=u$ (for $\theta=s$ the argument is similar) and call $\Lambda^{u}(t_{n})=\Lambda^{n}$. Observe that $\mathcal{S}=\{\Lambda^{n}\}_{n\in\mathcal{I}}$ is the sequence commented in the introduction of the section.
@@ 3.6. Subhorseshoes and connection by periodic orbits
In this subsection, we associate to every term of the sequence $\mathcal{S}$ a periodic orbit with the property that if $\Lambda^{n}$ and $\Lambda^{m}$ are associated with the same periodic orbit then they connect before $\max\{t_{n},t_{m}\}$.
In order to do that, given some $n$, remember the construction of $\Lambda^{n}$ given by proposition 3.12. A close inspection of the proof of that proposition shows that for some maximal invariant set, said $M^{n}$, that contains $\Lambda_{t_{n}}$ we took the subhorseshoe with maximal Hausdorff dimension $\Lambda_{0}^{n}\subset M^{n}$ and then applied proposition 3.5 in order to obtain the subhorseshoe $\Lambda^{n}$ with
@@
Next, if $D_{u}(M^{n})=D_{u}(\Lambda_{2}^{n})$ where $\Lambda_{2}^{n}\subset M^{n}$ is a subhorseshoe of $\Lambda$, then as $\Lambda_{0}^{n}$ has maximal dimension, it follows that either $D_{u}(\Lambda_{2}^{n})\leq D_{u}(\Lambda_{0}^{n})$ or $D_{s}(\Lambda_{2}^{n})\leq D_{s}(\Lambda_{0}^{n})$. In the first case
$D_{u}(\Lambda_{t_{n}})\leq D_{u}(M^{n})=D_{u}(\Lambda_{2}^{n})\leq D_{u}(\Lambda_{0}^{n})\leq\frac{r_{2}}{r_{1}}D_{u}(\Lambda_{0}^{n})$
and in the second, 3.11 let us conclude that
$D_{u}(\Lambda_{t_{n}})\leq D_{u}(M^{n})=D_{u}(\Lambda_{2}^{n})\leq r_{2}D_{s}(\Lambda_{2}^{n})\leq r_{2}D_{s}(\Lambda_{0}^{n})\leq\frac{r_{2}}{r_{1}}D_{u}(\Lambda_{0}^{n})$
@@
Now, take $r_{0}$ big enough such that $2^{2023}<N_{u}(\Lambda_{t_{0}},r_{0})$ and
@@
We set $\mathcal{B}_{0}=\mathcal{C}_{u}(\Lambda_{t_{0}},r_{0})$, $N_{0}=N_{u}(\Lambda_{t_{0}},r_{0})$ and for $n\in\mathcal{I}$, $M\in\mathbb{N}$ define the set
@@
[…]
Recall that the sizes of the intervals $I^{u}(\alpha)$ behave essentially submultiplicatively due the bounded distortion property of $\psi_{u}$ (equation 2.1) so that, one has
@@
$>$ $1.001\displaystyle 0.999^{4}/$ 1.001
$>$ $991/1000$
Then we have proved the next result
@@
Given $n\in\mathbb{N}$ and $M_{n}$ large
@@
with $\beta_{i}^{(p)}\in\mathcal{B}_{0}$ for all $41\leq i\leq j,\ p=3,$ 4 such that $\sup I^{s}((\beta_{j}^{(3)})^{T})<\inf I^{s}(\beta_{j}^{T})<\sup I^{s}(\beta_{j}^{T})<\inf I^{s}((\beta_{j}^{(4)})^{T})$, i.e., the interval $I^{s}(\beta_{j}^{T})$ is located between $I^{s}((\beta_{j}^{(3)})^{T})$ and $I^{s}((\beta_{j}^{(4)})^{T})$.
@@
The bounded distortion property (equation 2.2) let us fix $J\in\mathbb{N}$ big enough such that for $\beta_{1}\beta_{2}\dots\beta_{J}$ and $\beta_{J+1}\beta_{J+2}$ admissible with $\beta_{1},\beta_{2},\dots,\beta_{J},\beta_{J+1},\beta_{J+2}\in\mathcal{B}_{0}=\mathcal{C}_{u}(\Lambda_{t_{0}},r_{0})$
@@
For $N_{n}$ big enough, there exists $\beta_{n}\in\mathcal{B}_{5N_{n}k}(\Lambda^{n})$ such that the number of $5N_{n}k$-good positions of $\beta_{n}$ is greater or equal than $1049N_{n}k/$ 10.
@@
Since $\lvert\mathcal{B}_{5N_{n}k}(\Lambda^{n})\rvert>N_{0}^{991N_{n}k/200}$ (by lemma 3.15) and $2^{1+101N_{n}k/20}\cdot N_{0}^{99N_{n}k/20}<N_{0}^{991N_{n}k/200}$ (from our choices of $r_{0}$, $N_{0}$ large), we have that there exists some $\beta_{n}\in\mathcal{B}_{5N_{n}k}(\Lambda^{n})$ with less than $N_{n}k/10$, $5N_{n}k$-bad positions. That is, with at least $105N_{n}k-N_{n}k/10=49N_{n}k/$ 10 good positions. ∎
Given $n\in\mathcal{I}$ take $N_{n}$ big enough as in the lemma 3.17 and such that for two elements $x,y\in\Lambda$ if their kneading sequences coincide in the central block (centered at the zero position) of size $12N_{n}+$ 1 then $\lvert f(x)-f(y)\rvert<\eta_{n}/2$.
@@
If $\beta_{n}=\beta^{n}_{1}\beta^{n}_{2}\dots\beta^{n}_{5N_{n}k}$ with $\beta^{n}_{r}\in\mathcal{B}_{0}$ for $i=1,\dots,5N_{n}k$ is as in the previous lemma and for some $1<i<j<5N_{n}k$, the positions $i-1,i,j,j+1$ are $5N_{n}k$-good positions of $\beta_{n}$ and $j-i\geq J$. Then for each $i\leq s\leq j$ and $\bar{n}\in P(\beta^{n}_{s})$ if $\eta=\beta^{n}_{i-1}\beta^{n}_{i}\dots\beta^{n}_{j}\beta^{n}_{j+1}$ and $x\in R(\eta;P(\eta,s;\bar{n}))\cap\Lambda$ we have $f(x)<t_{n}$.
@@
The arguments are similar to those of proposition 2.9 of [1]. Consider $\underline{\theta}^{(2)}\in\mathcal{A}^{\mathbb{N}}$ and $\underline{\theta}^{(1)}\in\mathcal{A}^{\mathbb{Z}^{-}}$ such that $\underline{\theta}^{(1)}\beta^{n}_{i-1};\beta^{n}_{i}\beta^{n}_{i+1}\dots\beta^{n}_{j-1}\beta^{n}_{j}\beta^{n}_{j+1}\underline{\theta}^{(2)}\in\Sigma_{\mathcal{B}}$. With this notation, our task is equivalent to show that
@@
for all $10\leq\ell\leq m_{1}+m+m_{2}-$ 1 where $\beta^{n}_{i}=a_{1}\dots a_{m_{1}}$, $\beta^{n}_{i+1}\dots\beta^{n}_{j-1}=b_{1}\dots b_{m}$ and $\beta^{n}_{j}=d_{1}\dots d_{m_{2}}$.
@@
for any $\underline{\theta}^{(4)}\in\mathcal{A}^{\mathbb{N}}$. By 3.18, it follows that
@@
Next, we note that if $r_{0}\in\mathbb{N}$ is sufficiently large, $c_{5}e^{c_{1}}.|I^{s}((\beta^{n}_{i-1})^{T})|<c_{4}/2$ In particular, we have that
@@
In summary, from 3.23, 3.24, and 3.25 we deduce that 3.20 holds, as we wanted to see. ∎
Consider $\beta_{n}=\beta^{n}_{1}\beta^{n}_{2}\dots\beta^{n}_{5N_{n}k}$ and divide its position set $I=\{1,2,\dots,5N_{n}k\}$ in positions packages of size $N_{n}k$. In the central package $I^{*}=\{2N_{n}k+1,2N_{n}k+2,\dots,3N_{n}k\}$, the number of $5N_{n}k$-bad positions is less than $105N_{n}k-49N_{n}k/10=N_{n}k/$ 10 and then subdividing that package now in $N_{n}$ package of positions of size $k$ we can find some package of size $k$ with less than $k/10$, $5N_{n}k$-bad positions, said
@@
because $53JN_{0}^{2}<(16/5)JN_{0}^{2}=2k/$ 5. In this way, we obtain positions such that
@@
Since for $1\leq r\leq 3N_{0}^{2}$ the number of possibilities for $(\beta^{n}_{j_{r}},\beta^{n}_{j_{r}+1})$ is at most $N_{0}^{2}$, we conclude that for some different $1\leq r_{1}(n),r_{2}(n)\leq 3N_{0}^{2}$ one has
@@
Next, we see that if for some $m,n\in\mathcal{I}$ we have $O(m)=O(n)$ then it is possible to go from $\Lambda^{m}$ to $\Lambda^{n}$ without leaving $\Lambda_{\max\{t_{n},t_{m}\}}$ and staying arbitrarily close of the orbit of the periodic point $p=\Pi^{-1}(\overline{O(m)})$ for times arbitrarily big.
@@
because $\Lambda^{m}\subset\Lambda_{t_{m}-\eta_{m}}$ and as $j_{r_{1}(m)}-j_{r_{0}(m)}>2N_{m}k-N_{m}k=N_{m}k$ we have that $\sigma^{l}(x_{R})$ coincides with $\sigma^{l}(\theta_{m}^{1};\beta_{m}\theta_{m}^{2})$ in the central block of size $12N_{m}+$ 1 centered at the zero position.

### v2

@@ 3.5. Sequences of subhorseshoes
In this subsection, we suppose existence of an infinite sequence of discontinuities of the map $L_{\varphi,f}$ in some closed subinterval of $I_{\varphi,f}$ that does not contain the first accumulation point of the Lagrange spectrum and then construct arbitrary large finite sequences of subhorseshoes with some specific properties. Observe that here is the first time when we use the hypothesis of the diffeomorphism being close to a conservative one.
[…]
Fix $\varphi\in\mathcal{U}^{*}$ with $HD(\Lambda)<1$, $f\in\mathcal{P}_{\varphi,\Lambda}$, some closed subinterval $I\subset I_{\varphi,f}$ that does not contain $c_{\varphi,f}$ and suppose we have an infinite sequence of discontinuities of $L_{\varphi,f}$ with $s\in I$ for every $s$ in the sequence. Then, as $L_{\varphi,f}(\min I)\leq L_{\varphi,f}(s)\leq HD(\Lambda_{s})$, by (2.9) and (2.10)
@@
Now, we will associate to each $n$ a pair of subhorseshoes of $\Lambda$. In fact, the two subhorseshoes $\Lambda^{s}(t_{n})$ and $\Lambda^{u}(t_{n})$ are given by Proposition 3.12 considering some $0<\epsilon_{n}<\min\{0.001,(t_{n+1}-t_{n})/2\}$ and they satisfy
@@
  • $\Lambda^{s}(t_{n})$ does not connect with $\Lambda^{u}(t_{n})$ before $t_{n}$,
  • $\Lambda^{s}(t_{n})$ connects with $\Lambda^{u}(t_{n})$ before $t_{n+1}$.
@@
We can take $\theta\in\{s,u\}$ such that given $N\in\mathbb{N}$ arbitrary, there exists a sequence $n_{1}<n_{2}<...<n_{N}$ of elements of $\mathcal{I}$ (where $\mathcal{I}$ is the index set of the sequence $\{t_{n}\}$) such that for $i,j\in\{1,...,N\}$ with $i\neq j$, $\Lambda^{\theta}(t_{n_{i}})$ and $\Lambda^{\theta}(t_{n_{j}})$ do not connect before $\max\{t_{n_{i}},t_{n_{j}}\}.$
@@
We do the proof when the index set of the sequence is $\mathcal{I}=\{n\in\mathbb{Z}:n\geq 1\}$, and the other case follows similarly.
In the first case take a maximal $r_{1}$-chain beginning with 1; that is, a $r_{1}$-chain $1=n_{1}<n_{2}<...<n_{r_{1}}$ such that for any $n>n_{r_{1}}$, $1=n_{1}<n_{2}<...<n_{r_{1}}<n$ is not a $(r_{1}+1)$-chain and then $\Lambda^{s}(t_{n_{r_{1}}})$ does not connect with $\Lambda^{s}(t_{n})$ before $t_{n}$. Next take a maximal $r_{2}$-chain beginning with $n_{r_{1}}+1$: $n_{r_{1}}+1=n^{(r_{1})}_{1}<n^{(r_{1})}_{2}<\dots<n^{(r_{1})}_{r_{2}}$ then, as before, for $n^{(r_{1})}_{r_{2}}<n$, $\Lambda^{s}(t_{n^{(r_{1})}_{r_{2}}})$ does not connect with $\Lambda^{s}(t_{n})$ before $t_{n}$. Now consider a maximal $r_{3}$-chain beginning with $n^{(r_{1})}_{r_{2}}+1$: $n^{(r_{1})}_{r_{2}}+1=n^{(r_{1},r_{2})}_{1}<n^{(r_{1},r_{2})}_{2}<\dots<n^{(r_{1},r_{2})}_{r_{3}}$ then for $n^{(r_{1},r_{2})}_{r_{3}}<n$, $\Lambda^{s}(t_{n^{(r_{1},r_{2})}_{r_{3}}})$ does not connect with $\Lambda^{s}(t_{n})$ before $t_{n}$.
@@
such that for $k_{1},k_{2}\geq 2$ with $k_{1}\neq k_{2}$, $\Lambda^{s}(t_{\tilde{n}_{k_{1}}})$ and $\Lambda^{s}(t_{\tilde{n}_{k_{2}}})$ does not connect before $\max\{t_{\tilde{n}_{k_{1}}},t_{\tilde{n}_{k_{2}}}\}$.
On the other hand, in the second case, take $r\in\mathbb{N}$ arbitrarily big and $n_{1}<n_{2}<...<n_{r}$ some $r$-chain, then we affirm that for $i,j\in\{1,...,r\}$ with $i\neq j$, $\Lambda^{u}(t_{n_{i}})$ and $\Lambda^{u}(t_{n_{j}})$ does not connect before $\max\{t_{n_{i}},t_{n_{j}}\}.$ In other case if for some $i_{0},j_{0}\in\{1,...,r\}$ with $i_{0}<j_{0}$, $\Lambda^{u}(t_{n_{i_{0}}})$ and $\Lambda^{u}(t_{n_{j_{0}}})$ connect before $t_{n_{j_{0}}}$ then as by Corollary 2.7, $\Lambda^{s}(t_{n_{j_{0}}})$ connect with $\Lambda^{s}(t_{n_{i_{0}}})$ before $t_{n_{j_{0}}}$ and as also $\Lambda^{s}(t_{n_{i_{0}}})$ connects with $\Lambda^{u}(t_{n_{i_{0}}})$ before $t_{n_{i_{0}}+1}$ (and then before $t_{n_{j_{0}}}$). Applying two times more that corollary we have that $\Lambda^{s}(t_{n_{j_{0}}})$ connects with $\Lambda^{u}(t_{n_{j_{0}}})$ before $t_{n_{j_{0}}}$ that is a contradiction. From this follows the result. ∎
Without loss of generality, we will suppose that in the previous proposition $\theta=u$ (for $\theta=s$ the argument is similar) and call $\Lambda^{u}(t_{n})=\Lambda^{n}$.
@@ 3.6. Subhorseshoes and connection by periodic orbits
In this subsection, we associate to every term of the sequence $\{\Lambda^{n}\}_{n\in\mathcal{I}}$ a periodic orbit with the property that if $\Lambda^{n}$ and $\Lambda^{m}$ are associated with the same periodic orbit then they connect before $\max\{t_{n},t_{m}\}$.
In order to do that, given some $n$, remember the construction of $\Lambda^{n}$ given by Proposition 3.12. A close inspection of the proof of that proposition shows that for some maximal invariant set, said $P^{n}$, that contains $\Lambda_{t_{n}}$ we took the subhorseshoe with maximal Hausdorff dimension $\Lambda_{0}^{n}\subset P^{n}$ and then applied Proposition 3.5 in order to obtain the subhorseshoe $\Lambda^{n}$ with
@@
Next, if $D_{u}(P^{n})=D_{u}(\Lambda_{2}^{n})$ where $\Lambda_{2}^{n}\subset P^{n}$ is a subhorseshoe of $\Lambda$, then as $\Lambda_{0}^{n}$ has maximal dimension, it follows that either $D_{u}(\Lambda_{2}^{n})\leq D_{u}(\Lambda_{0}^{n})$ or $D_{s}(\Lambda_{2}^{n})\leq D_{s}(\Lambda_{0}^{n})$. In the first case
$D_{u}(\Lambda_{t_{n}})\leq D_{u}(P^{n})=D_{u}(\Lambda_{2}^{n})\leq D_{u}(\Lambda_{0}^{n})\leq\frac{r_{2}}{r_{1}}D_{u}(\Lambda_{0}^{n})$
and in the second, (3.13) let us conclude that
$D_{u}(\Lambda_{t_{n}})\leq D_{u}(P^{n})=D_{u}(\Lambda_{2}^{n})\leq r_{2}D_{s}(\Lambda_{2}^{n})\leq r_{2}D_{s}(\Lambda_{0}^{n})\leq\frac{r_{2}}{r_{1}}D_{u}(\Lambda_{0}^{n})$
@@
Now, by (2.5), we can take $r_{0}$ big enough such that $2^{2023}<N_{u}(\Lambda_{t_{0}},r_{0})$ and
@@
Set $\mathcal{B}_{0}=\mathcal{C}_{u}(\Lambda_{t_{0}},r_{0})$, $N_{0}=N_{u}(\Lambda_{t_{0}},r_{0})$ and for $n\in\mathcal{I}$, $M\in\mathbb{N}$ define the set
@@
Before continuing, we introduce some notation. Consider $\beta=\beta_{k_{1}}\beta_{k_{2}}...\beta_{k_{\ell}}=a_{1}...a_{p}\in\mathcal{A}^{p},\ \beta_{k_{i}}\in\mathcal{B}_{0},\ 1\leq i\leq\ell$. We say that $n\in\{1,...,p\}$ is the n-th position of $\beta$. If $\beta_{k_{i}}\in\mathcal{A}^{n_{k_{i}}}$ we write $|\beta_{k_{i}}|=n_{k_{i}}$ for its length and $P(\beta_{k_{i}})=\{1,2,...,n_{k_{i}}\}$ for its set of positions as a word in the alphabet $\mathcal{A}$ and given $s\in P(\beta_{k_{i}})$ we call $P(\beta,k_{i};s)=n_{k_{1}}+...+n_{k_{i-1}}+s$ the position in $\beta$ of the position $s$ of $\beta_{k_{i}}$.
Recall that the sizes of the intervals $I^{u}(\alpha)$ behave essentially submultiplicatively due the bounded distortion property of $\psi_{u}$ (equation (2.1)) so that, one has
@@
$>$ $0.999^{4}/1.001$
$>$ $991/1000.$
Thus we have proved the next result
@@
Given $n\in\mathcal{I}$ and $M_{n}$ large
@@
with $\beta_{i}^{(p)}\in\mathcal{B}_{0}$ for all $1\leq i\leq j,\ p=3,4$ such that $\sup I^{s}((\beta_{j}^{(3)})^{T})<\inf I^{s}(\beta_{j}^{T})<\sup I^{s}(\beta_{j}^{T})<\inf I^{s}((\beta_{j}^{(4)})^{T})$, i.e., the interval $I^{s}(\beta_{j}^{T})$ is located between $I^{s}((\beta_{j}^{(3)})^{T})$ and $I^{s}((\beta_{j}^{(4)})^{T})$.
@@
The bounded distortion property (Equation (2.2)) let us fix $J\in\mathbb{N}$ big enough such that for $\beta_{1}\beta_{2}\dots\beta_{J}$ and $\beta_{J+1}\beta_{J+2}$ admissible with $\beta_{1},\beta_{2},\dots,\beta_{J},\beta_{J+1},\beta_{J+2}\in\mathcal{B}_{0}=\mathcal{C}_{u}(\Lambda_{t_{0}},r_{0})$
@@
For $N_{n}$ big enough, there exists $\beta_{n}\in\mathcal{B}_{5N_{n}k}(\Lambda^{n})$ such that the number of $5N_{n}k$-good positions of $\beta_{n}$ is greater or equal than $49N_{n}k/10$.
@@
Since $\lvert\mathcal{B}_{5N_{n}k}(\Lambda^{n})\rvert>N_{0}^{991N_{n}k/200}$ (by Lemma 3.15) and $2^{1+101N_{n}k/20}\cdot N_{0}^{99N_{n}k/20}<N_{0}^{991N_{n}k/200}$ (from our choices of $r_{0}$, $N_{0}$ large), we have that there exists some $\beta_{n}\in\mathcal{B}_{5N_{n}k}(\Lambda^{n})$ with less than $N_{n}k/10$, $5N_{n}k$-bad positions. That is, with at least $5N_{n}k-N_{n}k/10=49N_{n}k/10$ good positions. ∎
Given $n\in\mathcal{I}$ take $N_{n}$ big enough as in Lemma 3.17 and such that for two elements $x,y\in\Lambda$ if their kneading sequences coincide in the central block (centered at the zero position) of size $2N_{n}+1$ then $\lvert f(x)-f(y)\rvert<\eta_{n}/2$ where $\Lambda^{s}(t_{n})\cup\Lambda^{u}(t_{n})\subset\Lambda_{t_{n}-{\eta_{n}}}$.
@@
If $\beta_{n}=\beta^{n}_{1}\beta^{n}_{2}\dots\beta^{n}_{5N_{n}k}$ with $\beta^{n}_{r}\in\mathcal{B}_{0}$ for $i=1,\dots,5N_{n}k$ is as in the previous lemma and for some $1<i<j<5N_{n}k$, the positions i1,i,j,j+1i-1,i,j,j+1 are $5N_{n}k$-good positions of $\beta_{n}$ and $j-i\geq J$. Then for each $i\leq s\leq j$ and $\bar{n}\in P(\beta^{n}_{s})$ if $\eta=\beta^{n}_{i-1}\beta^{n}_{i}\dots\beta^{n}_{j}\beta^{n}_{j+1}$ and $x\in R(\eta;P(\eta,s;\bar{n}))\cap\Lambda$ we have $f(x)<t_{n}$.
@@
The arguments are similar to those of [1, Proposition 2.9]. Consider $\underline{\theta}^{(2)}\in\mathcal{A}^{\mathbb{N}}$ and $\underline{\theta}^{(1)}\in\mathcal{A}^{\mathbb{Z}^{-}}$ such that $\underline{\theta}^{(1)}\beta^{n}_{i-1};\beta^{n}_{i}\beta^{n}_{i+1}\dots\beta^{n}_{j-1}\beta^{n}_{j}\beta^{n}_{j+1}\underline{\theta}^{(2)}\in\Sigma_{\mathcal{B}}$. With this notation, our task is equivalent to show that
@@
for all $0\leq\ell\leq m_{1}+m+m_{2}-1$ where $\beta^{n}_{i}=a_{1}\dots a_{m_{1}}$, $\beta^{n}_{i+1}\dots\beta^{n}_{j-1}=b_{1}\dots b_{m}$ and $\beta^{n}_{j}=d_{1}\dots d_{m_{2}}$.
@@
for any $\underline{\theta}^{(4)}\in\mathcal{A}^{\mathbb{N}}$. By (3.20), it follows that
@@
Next, we note that if $r_{0}\in\mathbb{N}$ is sufficiently large, $c_{5}e^{c_{1}}.|I^{s}((\beta^{n}_{i-1})^{T})|<c_{4}/2$ . In particular, we have that
@@
In summary, from (3.25), (3.26), and (3.27) we deduce that (3.22) holds, as we wanted to see. ∎
Consider $\beta_{n}=\beta^{n}_{1}\beta^{n}_{2}\dots\beta^{n}_{5N_{n}k}$ and divide its position set $I=\{1,2,\dots,5N_{n}k\}$ in positions packages of size $N_{n}k$. In the central package $I^{*}=\{2N_{n}k+1,2N_{n}k+2,\dots,3N_{n}k\}$, the number of $5N_{n}k$-bad positions is less than $5N_{n}k-49N_{n}k/10=N_{n}k/10$ and then subdividing that package now in $N_{n}$ package of positions of size $k$ we can find some package of size $k$ with less than $k/10$, $5N_{n}k$-bad positions, said
@@
because $3JN_{0}^{2}<(16/5)JN_{0}^{2}=2k/5$. In this way, we obtain positions such that
@@
Since for $1\leq r\leq 3N_{0}^{2}$ the number of options for $(\beta^{n}_{j_{r}},\beta^{n}_{j_{r}+1})$ is at most $N_{0}^{2}$, we conclude that for some different $1\leq r_{1}(n),r_{2}(n)\leq 3N_{0}^{2}$ one has
@@
Next, we see that if for some $m,n\in\mathcal{I}$ we have $O(m)=O(n)$ then it is possible to go from $\Lambda^{m}$ to $\Lambda^{n}$ without leaving $\Lambda_{\max\{t_{n},t_{m}\}}$ and staying arbitrarily close of the orbit of the periodic point $p=\Pi^{-1}(\overline{O(m)})$ for times arbitrarily big. More precisely, we have the following result
@@
because $\Lambda^{m}\subset\Lambda_{t_{m}-\eta_{m}}$ and as $j_{r_{1}(m)}-j_{r_{0}(m)}>2N_{m}k-N_{m}k=N_{m}k$ we have that $\sigma^{l}(x_{R})$ coincides with $\sigma^{l}(\theta_{m}^{1};\beta_{m}\theta_{m}^{2})$ in the central block of size $2N_{m}+1$ centered at the zero position.